In hospital, there are 12 nurses and 4 doctors. If a committee contains 6 persons, what is the probability that two doctors in this committee?
Answer by stanbon(75887)

In hospital, there are 12 nurses and 4 doctors. If a committee contains 6 persons, what is the probability that two doctors in this committee?
-----
# of ways to pick 2 doctors:: 4C2 = 6
# of ways to pick 4 nurses:: 12C4 = (12*11*10*9)/(1*2*3*4) = 99*5 = 495
----
# of committies with 2 docs and nurses = 6*495
# of random committies with 6 persons = 16C6
-------------
P(condition) = (6*495)/16C6 = (6*495)/8008 = 0.3709
Cheers,
Stan H.
-------------